A bottle of soda contains 591 mL of liquid. What is the volume in m3?


1
Expert's answer
2021-03-08T05:58:28-0500

1 mL = 10-6 m3

591 mL = 591×10-6 m3

591 mL = 5.91 ×10-4 m3